# Debiased Nonparametric Regression for Statistical Inference and Distributionally Robustness

### v2

@@ 3.1 Bias and variance decomposition
First, we investigate the bias and variance of $\widetilde{f}_{n}$. That is, we aim to analyze the bias and variance of $\widetilde{b}_{n}(x_{0})$, denoted as
@@
$\mathrm{Variance}(x_{0})$ $\coloneqq{\mathbb{E}}\left[\Big{(}\widetilde{f}_{n}(x_{0})-{\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\right]\Big{)}^{2}\right].$
Let us introduce the population values of ^$\widehat{{\mathcal{B}}}_{n,x}$ and $\widehat{s}_{n,x}$ as
@@
For this value, the covariate distribution, and the smoothness of $f_{0}$ and $\widehat{f}_{n}$, we make the following assumptions.
Assumption 3.2.
Let $s,L,C>0$ be constants independent of $f_{0}$ and $n$. Let $h_{n}$ be the bandwidth of the local polynomial estimator. For a point $x_{0}\in{\mathcal{X}}$ of interest, the following holds:
  • $f_{0}-\widehat{f}$ belongs to the Hölder class $\Sigma(s,L)$.
  • The eigenvalues of ^$\widehat{{\mathcal{B}}}_{n}(x)$ is bounded above and below away from zero.
  • There exists a constant $C>0$ such that $\int\mathbbm{1}\big{[}\|x-x_{0}\|\leq h\big{]}{\mathrm{d}}P(x)\leq Ch^{d}$,
where $P(x)$ denotes a distribution of $x$.
@@
Let $s,L,C>0$ be constants independent of $f_{0}$ and $n$. For the values, suppose that Assumptions 3.2 holds. Let $h_{n}$ be the bandwidth of the local polynomial estimator. Then, for any $\varepsilon>0$ and for all $x_{0}\in[0,1]$, there exits $n_{0}>0$ such that for all $n\geq n_{0}$, with probability $1-\varepsilon$, it holds that
$\left|\mathrm{Bias}(x_{0})\right|\leq C_{1}h^{s}_{n},$
$\mathrm{Variance}(x_{0})\leq\frac{C_{2}}{nh^{d}_{n}}+C_{3}h^{2s}_{n},$
where $C_{1},C_{2},C_{3}>0$ are constants independent of $f_{0}$ and $n$.
The smoothness condition on $f_{0}-\widehat{f}$ can be satisfied in several ways. We discuss this point in Section 4.
3.2 Pointwise convergence of the mean squared errors
Theorem 3.3 about the convergences of bias and variance terms directly indicate the convergence of the mean squared errors. This is because we have
${\mathbb{E}}\left[\left(\widehat{f}_{n}(x_{0})-f_{0}(x_{0})\right)^{2}\right]=\mathrm{Bias}(x_{0})^{2}+\mathrm{Variance}(x_{0})$
To minimize the mean squared error, we set the bandwidth as
$h_{n}=\alpha n^{-\frac{1}{2s+d}},$
where $\alpha$ is a constant independent of $f_{0}$ and $n$, and recall that $s$ is the parameter of smoothness and $d$ is the dimension of $X$. Then, we state the pointwise convergence of the mean squared errors as follows:
Theorem 3.4 (Pointwise mean squared error convergence).
Suppose that the conditions in Theorem 3.3 holds. Let $h_{n}=\alpha n^{-\frac{1}{2s+d}}$. Then, for all $x_{0}\in[0,1]$, it holds that
$\limsup_{n\to\infty}\sup_{f\in\Sigma(s,L)}\sup_{x_{0}\in{\mathcal{X}}}{\mathbb{E}}_{f}\left[\psi^{-2}_{n}\Big{|}\widetilde{f}_{n}(x_{0})-f(x_{0})\Big{|}^{2}\right]\leq C<\infty,$
where $\psi_{n}\coloneqq n^{-\frac{s}{2s+d}}$ is the rate of convergence and $C$ is a constant depending only on $s$, $L$, $\lambda_{0}$, $a_{0}$, $\sigma^{2}_{\max}$, $K_{\max}$, and $s$.
Theorem 3.4 can be extend to the following mean squared error convergence over the distribution of $X$.
Corollary 3.5 (Mean squared errors over the distribution of $X$).
Suppose that the conditions in Theorem 3.4 holds. Then, it holds that
$\limsup_{n\to\infty}\sup_{f\in\Sigma(s,L)}{\mathbb{E}}_{f}\left[\psi^{-2}_{n}\Big{\|}\widetilde{f}_{n}-f\Big{\|}^{2}_{2}\right]\leq C<\infty,$
where $\psi_{n}\coloneqq n^{-\frac{s}{2s+d}}$ is the rate of convergence and $C$ is a constant depending only on $s$, $L$, $\lambda_{0}$, $a_{0}$, $\sigma^{2}_{\max}$, $K_{\max}$, and $s$.
@@ 3.3 Pointwise asymptotic normality
Suppose one of the upper bounds in Theorem 3.4 holds. Then, it tholds that
$\sqrt{nh^{d}_{n}}\Big{(}\widetilde{f}(x_{0})-f_{0}(x_{0})\Big{)}\xrightarrow{{\mathrm{d}}}{\mathcal{N}}(0,V(x_{0})),$
where $V(x_{0})\coloneqq nh^{d}_{n}\mathrm{Variance}(x_{0})\leq C_{2}+o(h_{n})$.
The asymptotic normality result facilitates statistical inference. However, constructing a confidence interval is challenging due to the unknown parameter $s$ involved in $h_{n}$. To address this issue, the concept of honest confidence intervals has been developed. For example, the bootstrap method can be employed to obtain valid confidence intervals, even when certain parameters remain unspecified.
@@ 3.4 Uniform convergence
Lastly, we establish the uniform (or sup-norm) convergence of $\widetilde{f}(x_{0})$. Uniform convergence bounds the supremum of the estimation error uniformly over $X\in{\mathcal{X}}$, without relying on the distribution of $X$, whereas mean squared convergence depends on the distribution of $X$. This theoretical guarantee is particularly useful for addressing distribution shift problems, such as covariate shift (Shimodaira, 2000). In covariate shift scenarios, the distribution of $X$ differs between the training and test data. While mean squared convergence may be affected by such a shift, uniform convergence remains unaffected, ensuring robust performance for both training and test data. Consequently, uniform convergence is critical in discussions involving distribution shifts.
@@
Suppose that $f_{0}$ belongs to a Hölder class $\Sigma(s,L)$ on $\mathcal{X}$, where $s>0$ and $L>0$. Let $\widehat{f}_{n}$ be the $\ell$) estimator of order $\ell=\lfloor s\rfloor$ with bandwidth
$h_{n}\coloneqq\alpha\left(\frac{\log(n)}{n}\right)^{\frac{1}{2s+d}}$
for some $\alpha>0$. Suppose the following:
  • Assumption 3.2 holds.
  • $\varepsilon_{i}$ is a sub-Gaussian random variable satisfying
    ${\mathbb{E}}\big{[}\exp\big{(}\lambda\varepsilon_{i}\big{)}\big{]}\leq\exp\big{(}K^{2}\lambda^{2}\big{)},\quad\forall\lambda\in{\mathbb{R}},$
@@
  • $K$ is a Lipschitz kernel: $K\in\Sigma(1,L_{K})$ on $\cal{X}$ with $0<L_{K}<\infty$.
Then, there exists a constant $C<\infty$ such that
$\limsup_{n\to\infty}\sup_{f\in\Sigma(s,L)}{\mathbb{E}}_{f}\big{[}\psi^{-2}_{n}\big{\|}\widetilde{f}_{n}-f\big{\|}^{2}_{\infty}\big{]}\leq C,$
where
$\psi_{n}\coloneqq\left(\frac{\log(n)}{n}\right)^{\frac{s}{2s+d}}.$
Uniform convergence provides robustness against distributional shifts but has not been proven for many nonparametric regression estimators produced by machine learning methods. This limitation arises partly from their data-adaptive nature and the reliance on empirical process techniques within the empirical risk minimization framework (van de Geer, 2009). Using empirical processes and model complexity, estimation error can be evaluated in a general manner, avoiding significant dependence on specific algorithms. However, such approaches often focus on population risk, such as mean squared error, without addressing how these results translate into uniform convergence guarantees.
For example, Schmidt-Hieber & Zamolodtchikov (2024) tackle this issue by proposing an estimator based on neural networks, demonstrating restricted uniform convergence in the case of one-dimensional covariates. In contrast, our proposed method is model-free and accommodates multiple-dimensional covariates, offering broader applicability and robustness.
@@ 3.5 Double robustness
Our proposed debiased estimator possesses the property of double robustness; that is, if either $\widehat{b}_{n}$ or $\widehat{f}_{n}$ is consistent, the resulting debiased estimator $\widetilde{f}_{n}$ is also consistent.
Recall that the debiased estimator is defined as
$\widetilde{f}_{n}(x_{0})=\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}Y_{i}-\widehat{f}_{n}(X_{i})\Big{)}\widehat{w}_{h}(X_{i},x_{0})+\widehat{f}_{n}(x_{0}).$
Suppose that $\widehat{f}_{n}(x_{0})$ converges in probability to $f^{\dagger}(x_{0})$ and $\widehat{b}_{n}(x_{0})$ converges in probability to $b^{\dagger}(x_{0})$. We examine the behavior of $\widetilde{f}_{n}(x_{0})$ under two cases: $f^{\dagger}(x_{0})=f_{0}(x_{0})$ and $b^{\dagger}(x_{0})=f_{0}(x_{0})-f^{\dagger}(x_{0})$.
First, consider the case where $f^{\dagger}(x_{0})=f_{0}(x_{0})$ and $b^{\dagger}(x_{0})\neq f_{0}(x_{0})-f^{\dagger}(x_{0})$. Then, it follows that
@@
where $w^{\dagger}(X_{i},x_{0})$ is a function depending only on $X_{i}$ and $x_{0}$. Since the first term, $\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}Y_{i}-f_{0}(X_{i})\Big{)}w^{\dagger}(X_{i},x_{0})$, converges in probability to zero, $\widetilde{f}_{n}(x_{0})$ converges in probability to $f_{0}(x_{0})$.
Next, consider the case where $f^{\dagger}(x_{0})\neq f_{0}(x_{0})$ and $b^{\dagger}(x_{0})=f_{0}(x_{0})-f^{\dagger}(x_{0})$. Then, it holds that
$\widetilde{f}_{n}(x_{0})=f_{0}(x_{0})-f^{\dagger}(x_{0})+f^{\dagger}(x_{0})+o_{p}(1)=f_{0}(x_{0})+o_{p}(1).$
Thus, in either case, $\widetilde{f}_{n}(x_{0})$ converges in probability to $f_{0}(x_{0})$. We summarize this property in the following theorem.
@@
If either $\widehat{b}_{n}$ or $\widehat{f}_{n}$ is consistent, then it holds that
$\widetilde{f}_{n}(x_{0})\xrightarrow{{\mathrm{p}}}f_{0}(x_{0})\qquad\mathrm{as}\ \ n\to\infty.$
@@ 4 Smoothness is all you need
As demonstrated above, our method requires only the smoothness of $\widehat{f}-f_{0}$. This assumption is satisfied when $f_{0}$ and $\widehat{f}$ are individually smooth, but it also holds in cases where $f_{0}$ and $\widehat{f}$ are not smooth, provided that their difference is smooth. Neural networks belong to this class under the analysis of (Schmidt-Hieber, 2020).
Remarkably, our results require only the smoothness of the first-stage estimator (the difference between $\widehat{f}_{n}$ and $f_{0}$). This assumption significantly broadens the applicability of our method. Even when $f_{0}$ and $\widehat{f}_{n}$ are not smooth, our method can still guarantee mean squared convergence, asymptotic normality, and uniform convergence. For instance, Imaizumi & Fukumizu (2019) discusses the convergence of nonparametric estimators using neural networks when $f_{0}$ belongs to a piecewise Hölder class. Under such a case, the liner estimators, including locally polynomial regression, are suboptimal. However, even under the case, the difference $\widehat{f}-f_{0}$ can be smooth, and our smoothness condition still holds.

### v3

@@ 4.2 Bias and variance decomposition
We begin by analyzing the bias and variance of $\widetilde{f}_{n}$, defined as
@@
$\mathrm{Variance}(x_{0})$ $\coloneqq{\mathbb{E}}\left[\left(\widetilde{f}_{n}(x_{0})-{\mathbb{E}}[\widetilde{f}_{n}(x_{0})]\right)^{2}\right].$
The bias term quantifies how closely the estimator $\widetilde{f}_{n}$ centers around the true regression function $f_{0}$ at $x_{0}$, while the variance term measures the fluctuation of $\widetilde{f}_{n}$ around its own mean. Balancing bias and variance is a fundamental concern in statistical estimation. Our approach, which employs local polynomial smoothing in the second stage, aims to systematically reduce the bias introduced by the first-stage estimator while controlling variance through appropriate bandwidth selection.
To facilitate our analysis, we introduce the population version of ^$\widehat{{\mathcal{B}}}_{n}(x_{0})$:
@@
We then establish bounds for the bias and variance terms as follows:
@@
Let $s,L,C,C_{1},C_{2},C_{3}>0$ be constants independent of $f_{0}$ and $n$. Let $h_{n}$ be the bandwidth of the local polynomial estimator. For every $x\in{\mathcal{X}}$, the following hold:
  • The eigenvalues of ~$\widetilde{{\mathcal{B}}}_{n}(x)$ are uniformly bounded above and below away from zero.
  • There exists a constant $C>0$ such that
    $\int\mathbbm{1}\big{[}|x-x_{0}|\leq h_{n}\big{]}{\mathrm{d}}P(x)\leq Ch_{n},$
    where $P(x)$ denotes the distribution of $x$.
If $f_{0}-\widehat{f}_{n}$ belongs to the Hölder class $\Sigma(s,L)$ almost surely as $n\to\infty$, then for any $\varepsilon>0$ and for all $x_{0}\in{\mathcal{X}}$, there exists $n_{0}>0$ such that for all $n\geq n_{0}$, with probability at least $1-\varepsilon$, it holds that
$\bigl{|}\mathrm{Bias}(x_{0})\bigr{|}$ $\leq C_{1}h^{s}_{n},$
$\mathrm{Variance}(x_{0})$ $\leq\frac{C_{2}}{nh_{n}}+C_{3}h^{2s}_{n}.$
The smoothness condition on $f_{0}-\widehat{f}_{n}$ can be satisfied in various ways; see Section 5 for further discussion. The above result implies that as long as the smoothness of $f_{0}-\widehat{f}_{n}$ is well-controlled and the bandwidth $h_{n}$ is chosen appropriately to shrink with $n$, both bias and variance remain bounded and ultimately vanish as $n$ grows.
4.3 Pointwise convergence of the MSE
From Theorem 4.2, the bounds on bias and variance imply the convergence of the MSE, noting that
${\mathbb{E}}\left[\left(\widehat{f}_{n}(x_{0})-f_{0}(x_{0})\right)^{2}\right]=\mathrm{Bias}(x_{0})^{2}+\mathrm{Variance}(x_{0}).$
To minimize the MSE, we set
$h_{n}=\alpha n^{-\frac{1}{2s+1}},$
where $\alpha$ is a constant independent of $f_{0}$ and $n$, $s$ is the smoothness parameter, and $d$ is the dimension of $X$. We then establish the pointwise MSE convergence result.
Theorem 4.3 (Pointwise MSE convergence).
Let $s,L,C,\alpha>0$ be constants independent of $f_{0}$ and $n$. Let
$h_{n}=\alpha n^{-\frac{1}{2s+1}}$
be the bandwidth of the local polynomial estimator. For every $x_{0}\in{\mathcal{X}}$, the following hold:
  • The eigenvalues of ~$\widetilde{{\mathcal{B}}}_{n}(x)$ are uniformly bounded above and below away from zero.
  • There exists a constant $C>0$ such that
    $\int\mathbbm{1}\big{[}|x-x_{0}|\leq h_{n}\big{]}{\mathrm{d}}P(x)\leq Ch_{n},$
    where $P(x)$ denotes the distribution of $x$.
If $f_{0}-\widehat{f}_{n}$ belongs to the Hölder class $\Sigma(s,L)$ almost surely as $n\to\infty$, then for all $x_{0}\in{\mathcal{X}}$, it holds that
${\mathbb{E}}\left[\left(\widetilde{f}_{n}(x_{0})-f_{0}(x_{0})\right)^{2}\right]=O\left(n^{-\frac{s}{2s+1}}\right),$
as $n\to\infty$.
Theorem 4.3 can be extended to establish the convergence of the MSE over the distribution of $X$.
Corollary 4.4 (MSE over the distribution of $X$).
Suppose the conditions in Theorem 4.3 hold. Then, we have
${\mathbb{E}}\left[\left(\widetilde{f}_{n}(X)-f_{0}(X)\right)^{2}\right]=O\left(n^{-\frac{2s}{2s+1}}\right),$
where the expectation is taken over both $\widetilde{f}_{n}$ and $X$, as $n\to\infty$.
@@ 4.4 Pointwise asymptotic normality
Let $s,L,C>0$ be constants independent of $f_{0}$ and $n$. Let $h_{n}$ be the bandwidth of the local polynomial estimator such that $h_{n}\to 0$, $nh_{n}\to\infty$, and $nh^{2s+1}_{n}\to 0$ hold as $n\to\infty$. For every $x\in{\mathcal{X}}$, the following hold:
  • The eigenvalues of ~$\widetilde{{\mathcal{B}}}_{n}(x)$ are uniformly bounded above and below away from zero.
  • There exists a constant $C>0$ such that
    $\int\mathbbm{1}\big{[}|x-x_{0}|\leq h\big{]}{\mathrm{d}}P(x)\leq Ch_{n},$
    where $P(x)$ is the distribution of $x$.
If $f_{0}-\widehat{f}_{n}$ belongs to the Hölder class $\Sigma(s,L)$ almost surely as $n\to\infty$, then for every $x_{0}\in{\mathcal{X}}$, it holds that
$\sqrt{nh_{n}}\left(\widetilde{f}(x_{0})-f_{0}(x_{0})\right)\xrightarrow{{\mathrm{d}}}{\mathcal{N}}\big{(}0,V(x_{0})\big{)},$
as $n\to\infty$, where $V(x_{0})\geq 0$ is a constant depending on $x_{0}$ but independent of $n$.
Here, note that the following holds from the bias and variance decomposition:
$V(x_{0})\coloneqq nh_{n}\mathrm{Variance}(x_{0})\leq C_{2}+o(h_{n}).$
The asymptotic normality result facilitates statistical inference, such as constructing confidence intervals or hypothesis tests for $f_{0}(x_{0})$. In practice, knowing that $\sqrt{nh_{n}}(\widetilde{f}(x_{0})-f_{0}(x_{0}))$ converges in distribution to a normal allows for classical inferential techniques, provided that the variance $V(x_{0})$ can be consistently estimated. However, constructing a confidence interval can be challenging, because $s$ (the smoothness parameter) is often unknown and affects the choice of $h_{n}$. In real applications, data-driven methods, including cross-validation, can be employed to select $h_{n}$, and bootstrap or other resampling approaches can help in variance estimation.
@@ 4.5 Uniform convergence
Finally, we establish the uniform (sup-norm) convergence of $\widetilde{f}_{n}(x_{0})$. Unlike mean squared convergence, which depends on the distribution of $X$, uniform convergence bounds the supremum of the estimation error over all $x\in{\mathcal{X}}$. This property is particularly valuable for addressing the covariate shift problem (Shimodaira, 2000; Schmidt-Hieber & Zamolodtchikov, 2024), as uniform convergence remains unaffected by changes in the distribution of $X$ between training and test datasets.
@@
Let $s,L_{K},L,C,\alpha>0$ be constants independent of $f_{0}$ and $n$. Let
$h_{n}\coloneqq\alpha\left(\frac{\log(n)}{n}\right)^{\frac{1}{2s+1}}$
be the bandwidth of the local polynomial estimator. For every $x_{0}\in{\mathcal{X}}$, the following hold:
  • The eigenvalues of ~$\widetilde{{\mathcal{B}}}_{n}(x)$ are uniformly bounded above and below away from zero.
  • There exists a constant $C>0$ such that
    $\int\mathbbm{1}\big{[}|x-x_{0}|\leq h_{n}\big{]}{\mathrm{d}}P(x)\leq Ch_{n},$
    where $P(x)$ denotes the distribution of $x$.
  • The noise $\varepsilon_{i}$ is sub-Gaussian, satisfying
    ${\mathbb{E}}\left[\exp(\lambda\varepsilon_{i})\right]\leq\exp\left(K^{2}\lambda^{2}\right)\quad\forall\lambda\in{\mathbb{R}},$
@@
  • The kernel function $K$ is Lipschitz: $K\in\Sigma(1,L_{K})$ on ${\mathcal{X}}$.
If $f_{0}-\widehat{f}_{n}$ belongs to the Hölder class $\Sigma(s,L)$ almost surely as $n\to\infty$, then the following holds:
${\mathbb{E}}\left[\left\|\widetilde{f}_{n}-f_{0}\right\|^{2}_{\infty}\right]=O\left(\frac{\log(n)}{n}\right)^{\frac{2s}{2s+1}}.$
Although uniform convergence ensures robust performance under distribution shifts, it has not been established for many nonparametric regression estimators derived from modern machine learning methods. One reason for this gap is their data-adaptive nature and the reliance on empirical process techniques, which primarily focus on population risk measures whose expectation is taken over the covariate distribution rather than on sup-norm bounds. For example, Schmidt-Hieber & Zamolodtchikov (2024) analyzes a neural network-based estimator in a one-dimensional covariate setting, demonstrating restricted uniform optimality. In contrast, our approach is model-free and accommodates multi-dimensional covariates, offering broader applicability and greater robustness.
@@ 4.6 Double robustness
Our proposed debiased estimator also possesses the property of double robustness: if either $\widehat{b}_{n}$ or $\widehat{f}_{n}$ is consistent, then $\widetilde{f}_{n}$ remains consistent.
Recall that
$\widetilde{f}_{n}(x_{0})=\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\left(Y_{i}-\widehat{f}_{n}(X_{i})\right)\widehat{w}_{h}(X_{i},x_{0})+\widehat{f}_{n}(x_{0}).$
Suppose $\widehat{f}_{n}(x_{0})$ converges in probability to $f^{\dagger}(x_{0})$ and $\widehat{b}_{n}(x_{0})$ converges in probability to $b^{\dagger}(x_{0})$. We analyze two cases:
  • If $f^{\dagger}(x_{0})=f_{0}(x_{0})$ but $b^{\dagger}(x_{0})\neq f_{0}(x_{0})-f^{\dagger}(x_{0})$, then
@@
    where $w^{\dagger}(X_{i},x_{0})$ depends only on $X_{i}$ and $x_{0}$. Since the first term converges in probability to zero, it follows that $\widetilde{f}_{n}(x_{0})\xrightarrow{{\mathrm{p}}}f_{0}(x_{0})$.
  • If $f^{\dagger}(x_{0})\neq f_{0}(x_{0})$ but $b^{\dagger}(x_{0})=f_{0}(x_{0})-f^{\dagger}(x_{0})$, then
    $\widetilde{f}_{n}(x_{0})=\left(f_{0}(x_{0})-f^{\dagger}(x_{0})\right)+f^{\dagger}(x_{0})+o_{p}(1)=f_{0}(x_{0})+o_{p}(1).$
In both cases, $\widetilde{f}_{n}(x_{0})\xrightarrow{{\mathrm{p}}}f_{0}(x_{0})$. We summarize this result as follows.
@@
For any $x_{0}\in{\mathcal{X}}$, if either $\widehat{b}_{n}(x_{0})$ or $\widehat{f}_{n}(x_{0})$ is consistent, then
$\widetilde{f}_{n}(x_{0})\xrightarrow{{\mathrm{p}}}f_{0}(x_{0})\qquad\text{as }n\to\infty.$
Double robustness thus provides an additional safeguard against model misspecification or suboptimal performance of one of the estimators. In practice, this property is particularly valuable when the first-stage model $\widehat{f}_{n}$ fails to accurately capture the true function, but the second-stage residual model $\widehat{b}_{n}$ remains reliable—or vice versa.
@@ 5 Smoothness is all you need
As demonstrated above, our method relies solely on the smoothness of $\widehat{f}-f_{0}$. This assumption holds when both $f_{0}$ and $\widehat{f}$ are individually smooth, but it also remains valid in cases where neither $f_{0}$ nor $\widehat{f}$ is smooth, as long as their difference is smooth. Neural networks satisfy this condition under the analysis in Schmidt-Hieber (2020).
Notably, our results require only the smoothness of the first-stage estimator (i.e., the difference $\widehat{f}_{n}-f_{0}$). This assumption significantly broadens the applicability of our method. Even when $f_{0}$ and $\widehat{f}_{n}$ are not individually smooth, our approach can still ensure MSE convergence, asymptotic normality, and uniform convergence, provided that the difference is smooth.
From a different perspective, this flexibility allows one to use state-of-the-art or specialized methods in the first stage without compromising the theoretical guarantees of the final estimator, as long as the first-stage error exhibits a level of smoothness akin to the Hölder condition. This feature is particularly useful when the true regression function $f_{0}$ has a complex or partially known structure, yet the difference $f_{0}-\widehat{f}_{n}$ remains more regular due to design choices or inductive biases in $\widehat{f}_{n}$.